OpenStudy (anonymous):

Don bought his car for $20,350. It is expected to depreciate an average of 11% each year during the first 5 years. What will the approximate value of his car be in 5 years

OpenStudy (anonymous):

After each year, the value of the car is only 100%-11%=89% of its previous year's value After one year, the value of the car is 20,350*89/100=... Compute this calculation five times and you'll get the answer.

OpenStudy (anonymous):

From the question, the depreciating value of his car is 0.11 * $20,350 each year. So, for five years multipy that by 5, subtract this value from the original cost to get the current value... \[$20,350 - [5 * (0.11 \times $20,350)] = ?\]

OpenStudy (anonymous):

11,363 ?

OpenStudy (anonymous):

That's much quicker formula, instead of having to do that for five times... ;)

OpenStudy (anonymous):

Indeed, but i'm just afraid that the op won't understand the shortcut. I was trying to give him/her a hint. :) In fact, we can do this: 20,350*(89/100)^5 :)

OpenStudy (anonymous):

I think there's something wrong with that concept "(100-11)%"

OpenStudy (anonymous):

If you have to go by that formula, then it would be (100-11n) % each year, where n is in years. That would make more sense.

OpenStudy (anonymous):

Check this: 11% in five years time= 11% x 5=55% therefore 100%-55%=45% Then: 45/100 x #20,350= $9,157.5 cost of car in five years time.

OpenStudy (anonymous):

yes, exactly, @Perculiar , (100 - 5*11) % We're on the same page here... :)

OpenStudy (anonymous):

I think you have interpreted the question in a different way. :D I was taking the value of the car at the end of each year and assume that that value then becomes only 89% of it, whereas you take the initial value and assume that it is going to decrease by the SAME amount after every year.

OpenStudy (anonymous):

Yes, 89% the first year, second year has got to have another decrease.. you see that?

OpenStudy (anonymous):

@09Ti08 I used your formula hope my approach is not too fatal.

OpenStudy (anonymous):

@d3Xter yeah! Second yeah would be 89% - 11% = 78% and so on...

OpenStudy (anonymous):

So, @katrinamarie0803, the right formula to use is the one I gave you before \[$20,350 - [5 * (0.11 \times $20,350)] = ?\]

OpenStudy (anonymous):

Good luck :)